OpenStudy (meilendurcer):

how do you solve 6/b-1=9/7?

OpenStudy (anonymous):

answer is 42/16 which is equal to 21/8

OpenStudy (anonymous):

\[\large \frac{6}{b-1}=\frac{9}{7}\iff 9(b-1)=6\times 7\]

OpenStudy (anonymous):

pretty sure the above answer is wrong

OpenStudy (meilendurcer):

i would think so

OpenStudy (anonymous):

\[9b-9=42\\ 9b=51\\ b=\frac{51}{9}=\frac{17}{3}\]

OpenStudy (anonymous):

mathway.com